tan 100 . + tan 125 . + (tan 100 . )(tan 125 . ) is equal to A. 0 B. 1/2 C. -1 D. 1 EXPLAIN HOW?

tan 100. + tan 125.+ (tan 100. )(tan 125.) is equal to


A. 0   B. 1/2    C.  -1    D.   1


EXPLAIN HOW?

Grade:12th Pass

1 Answers

G Narayana Raju
12 years ago

tan225=tan(125+100),

tan(180+45)= (tan125+tan100)/(1-tan125tan100),

tan45=(tan125+tan100)/(1-tan125tan100),

1=(tan125+tan100)/(1-tan125tan100),

tan125+tan100=1-tan125tan100,

tan125+tan100+tan125tan100=1.